[Hallc_running] [New Logentry] Beam Update

hszumila at jlab.org hszumila at jlab.org
Fri Feb 9 09:15:02 EST 2018

Logentry Text:
We had pretty smooth running overnight with slightly lower current. We finished our long coincidence running on the C target. Beam was taken away this morning and will be off through sometime Sunday. The purpose is to retune the beam line. This morning they will be going in to fix the mis-wired quad. I will have a more concrete time line tomorrow. In the meantime, the rest of day shift today through Sunday Owl shift is cancelled. The first shift that can be expected to be available will the day shift Sunday. 

For now, the plan is to cold park the target during this time and start re-condensing on Sunday during the day shift. The target will be taken over and parked by the target group around noon today. 

Everything is still tentative, but it is possible that Monday will not be dedicated to beam studies although the rf recovery is still scheduled for Tuesday day shift. We will meet in 200C at 1:30 pm (second floor of the counting house) to discuss the plan going forward.

Link: https://bluejeans.com/543903672
Connecting directly from a room system?
1) Dial: or bjn.vc
2) Enter Meeting ID: 543 903 672
Just want to dial in on your phone?
1) Direct-dial with my iPhone or
+1.408.740.7256 (United States)
+1.888.240.2560 (US Toll Free)
+1.408.317.9253 (Alternate number)
(Global Numbers)
2) Enter Meeting ID: 543 903 672
3) Press #


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3529693
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180209/04b8cbbe/attachment.html>

More information about the Hallc_running mailing list